# MedFormer: Hierarchical Medical Vision Transformer with Content-Aware Dual Sparse Selection Attention

### v1

@@ Abstract
Medical image recognition serves as a key way to aid in clinical diagnosis, enabling more accurate and timely identification of diseases and abnormalities. Vision transformer-based approaches have proven effective in handling various medical recognition tasks. However, these methods encounter two primary challenges. First, they are often task-specific and architecture-tailored, limiting their general applicability. Second, they usually either adopt full attention to model long-range dependencies, resulting in high computational costs, or rely on handcrafted sparse attention, potentially leading to suboptimal performance. To tackle these issues, we present MedFormer, an efficient medical vision transformer with two key ideas. First, it employs a pyramid scaling structure as a versatile backbone for various medical image recognition tasks, including image classification and dense prediction tasks such as semantic segmentation and lesion detection. This structure facilitates hierarchical feature representation while reducing the computation load of feature maps, highly beneficial for boosting performance. Second, it introduces a novel Dual Sparse Selection Attention (DSSA) with content awareness to improve computational efficiency and robustness against noise while maintaining high performance. As the core building technique of MedFormer, DSSA is explicitly designed to attend to the most relevant content. In addition, a detailed theoretical analysis has been conducted, demonstrating that MedFormer has superior generality and efficiency in comparison to existing medical vision transformers. Extensive experiments on a variety of imaging modality datasets consistently show that MedFormer is highly effective in enhancing performance across all three above-mentioned medical image recognition tasks. The code is available on GitHub.
@@ 1 Introduction
Due to their ability to provide rapid identification of lesion types and accurate localization of affected regions, various visual recognition tasks related to medical images – such as classification, segmentation, and detection – are extremely important for medical auxiliary diagnosis [1], [2]. Recently, vision transformer-based approaches demonstrate their effectiveness in addressing these medical recognition tasks. However, due to the diversity and complexity of medical images across different imaging modalities (e.g., MRI, CT, and Optics) and varying image quality (e.g., signal-to-noise ratio and artifacts), most existing vision transformer architectures are typically tailored for task-specific scenarios, such as disease diagnosis [3], [4], [5], [6], anatomical structure and lesion segmentation [7], [8], [9], [10], [11], and abnormality detection [12], [13]. While effective in their respective applications, such task-specific designs may lead to unsatisfactory performance when applied to diverse medical image recognition tasks. This limitation underscores the need for a general-purpose transformer architecture with task-agnostic, modality-invariant, and quality-adaptive characteristics.
@@
The main technical contributions – hierarchical feature representations through a pyramid structure and the DSSA mechanism – endow MedFormer with a strong ability to effectively model long-range dependencies. As a result, the proposed MedFormer architecture serves as a robust general-purpose backbone without any additional bells and whistles. This makes it particularly appropriate for medical recognition tasks such as disease diagnosis, anatomical structure segmentation, and abnormality detection. To the best of our knowledge, this is the first work to attempt developing a general-purpose vision transformer backbone for the medical image recognition community. Overall, our main contributions can be summarized as follows:
@@
The rest of this paper is organized as follows. Section II delves into the prior related works and their limitations. Section III specifies our method, including MedFormer architecture and DSSA mechanism. Section IV showcases the experimental results and details, particularly the visualization of DSSA attention map. Section V summarizes our findings and indicates directions for further research.
@@ 3.1.3 Matrix Multiplication
To enable matrix multiplication between $\mathbf{A}^{P}$ and $\mathbf{V}^{gg}$, we introduce an additional dimension to the third axis of $\mathbf{A}^{P}$, resulting in a tensor shape of $S^{2}\times\frac{HW}{S^{2}}\times 1\times\frac{k_{1}HW}{S^{2}}$. Furthermore, to enrich local feature information, we incorporate a local context enhancement module, denoted as $\mathrm{LCE}(\cdot)$, which is constructed using a $5\times 5$ depth-wise convolution. This module further strengthens feature representation. The final output of the model is expressed as:
@@
Table 1: Variants of MedFormer for medical image classification. $C_{i}$: Channel dimension at stage $i$. $S$: the number of regions in feature maps. $k_{1}$: Select the $k_{1}$ most relevant regions. $\lambda$: proportionality coefficient used for the second sparse selection.
@@ 3.2 Complexity Analysis of DSSA
$=3HWC^{2}+S^{2}(S^{2}C+S^{2}+k_{1})+HW\frac{{k_{1}}HW}{S^{2}}C$
$\quad+S^{2}\frac{HW}{S^{2}}(\frac{{k_{1}}HW}{S^{2}}+k_{2})+HW{k_{2}}C$
$<3HWC^{2}+2(S^{2})^{2}C+\frac{2k(HW)^{2}}{S^{2}}(C+1).$
Next, we denote the right side of the inequality in Eq. 11 as $\mathrm{InEq}$, and proceed with a further analysis:
$\mathrm{InEq}$ $=3HWC^{2}+2(S^{2})^{2}C+\frac{2k(HW)^{2}}{S^{2}}(C+1)$ (12)
$\geq 3HWC^{2}+3\left(2(S^{2})^{2}C\cdot 2HW\frac{kHW}{S^{2}}(C+1)\right)^{\frac{1}{3}}$
$=3HWC^{2}+6(Ck)^{\frac{2}{3}}(HW)^{\frac{4}{3}}.$
Combining Eq. 11 and Eq. 12, it is evident that the FLOPs must be less than $\mathrm{InEq}$.
$\mathrm{FLOPs}<3HWC^{2}+6(Ck)^{\frac{2}{3}}(HW)^{\frac{4}{3}}.$ (13)
@@
Classification ISIC-2018 Classification[36][37] $224\times 224$ 7 10015 193 1512
ColonPath [38] $224\times 224$ 2 8007 2002 -
Brain Tumor [39] $224\times 224$ 4 5712 1311 -
Segmentation ISIC-2018 Segmentation [36][37] $256\times 256$ 2 1868 467 259
CVC-ClinicDB [40] $256\times 256$ 2 441 110 61
Synapse [41] $224\times 224$ 9 2212 1567 -
Detection Kvasir-Seg [42] $256\times 256$ 2 880 - 120
Brain Tumor Detection [43] $256\times 256$ 4 6930 - 1980
Table 3: Comparison of different backbones on three medical image classification datasets. * indicates that the model is initialized with weights pretrained on the ImageNet dataset.
Methods Params (M) FLOPs (G) Top-1 Acc (%)
@@
HiFuse-Tiny[27] 120.00 17.56 84.46 98.70 95.27
MIL-VT[21] 21.80 4.28 83.94 97.55 87.19
KAT [30] 20.08 3.01 87.98 98.22 96.23
MG-Trans [31] 22.12 3.38 87.68 98.46 97.02
ViT-S[14] 21.67 4.61 87.11 97.30 88.25
PVT-S[19] 23.98 3.83 88.60 97.30 90.47
Swin-T[15] 27.53 4.51 87.63 98.10 91.38
BiFormer-S[16] 25.03 4.49 86.01 98.35 96.34
MedFormer-T 3.20 0.58 85.49 97.29 91.56
MedFormer-S 12.63 2.18 88.08 98.10 95.73
MedFormer-B 28.30 4.79 89.64 98.72 98.02
MedFormer-T* 3.20 0.58 93.44 100.00 99.69
MedFormer-S* 12.63 2.18 94.30 100.00 99.85
MedFormer-B* 28.30 4.79 94.30 100.00 99.92
@@ 3.3.1 MedFormer Block
Utilizing DSSA, we construct a basic block called the MedFormer Block. This block is composed of three main components: a $3\times 3$ depth-wise convolution to implicitly encode relative positional information, a DSSA layer performs sparse selection both at region-level and pixel-level to capture critical features, and a two-layer MLP module with expansion ratio $e=3$. Each component incorporates residual connections to maintain the model’s learning capacity, and layer normalization is applied to stabilize the training process and improve the capability of generalization. The detailed design of the block is shown in Fig. 3.
@@ 3.3.2 Architecture
To produce hierarchical feature maps for enhancing performance in classification and various dense prediction tasks, we construct a hierarchical pyramid medical backbone network based on MedFormer Blocks, named MedFormer. The detailed design is illustrated in Fig. 3. Specifically, Stage 1 consists of a patch embedding layer and $N_{1}$ MedFormer Blocks. The patch embedding layer is composed of two $3\times 3$ convolutions, which transform the input image into a feature map with a resolution of $\frac{H}{4}\times\frac{W}{4}\times C$. To obtain a hierarchical representation, Stage 2 introduces a patch merging layer with a $3\times 3$ convolution, which reduces the number of tokens and sets the channel dimension to $2C$, resulting in a feature map size of $\frac{H}{8}\times\frac{W}{8}\times 2C$. Then, $N_{2}$ MedFormer Blocks are applied to extract effective features. After repeating the process twice, Stage 3 and Stage 4 generate feature maps with dimensions $\frac{H}{16}\times\frac{W}{16}\times 4C$ and $\frac{H}{32}\times\frac{W}{32}\times 8C$, respectively. Additionally, by adjusting the parameters of MedFormer, we proposed three different variants: Tiny, Small and Base. Table 1 presents the detailed parameter configurations of these MedFormer variants for classification tasks.
@@ 4.1 Datasets
This subsection introduces the datasets used for each task, with detailed information about each dataset provided in Table 2.
@@ 4.1.1 ISIC-2018 Classification
The classification dataset originates from the ISIC-2018 Challenge Task 3 and encompasses seven distinct categories of images. All images have been standardized to a uniform resolution of $600\times 450$ pixels. The entire dataset consists of 10,015 training images, 193 validation images, and 1,512 test images.
@@ 4.1.2 ColonPath
This is a binary classification dataset from Ruijin Hospital in China, designed to detect malignant cell regions in tissue slides. These images are derived from colonoscopy results of 396 patients. We randomly split the dataset into 8,007 training images and 2,002 validation images.
@@ 4.1.3 Brain Tumor
This is an MRI image dataset for four-class brain tumor classification, consisting of a total of 7,023 images, including 5,712 training images and 1,311 validation images.
@@ 4.1.4 ISIC-2018 Segmentation
This segmentation dataset is collocated from the ISIC-2018 Challenge Task 1 and contains 2,594 images with varying resolutions. We randomly split the dataset into 1,868 training images, 467 validation images and 259 testing images.
@@ 4.1.5 CVC-ClinicDB
CVC-ClinicDB is a widely used polyp segmentation dataset consisting of 612 images. For our segmentation experiments, 441 images are randomly selected for training, 110 for validation, and 61 for testing.
@@ 4.1.6 Synapse
This is a multi-organ segmentation dataset consists of 30 abdominal CT scans, with each CT volume containing between 85 and 198 slices of $512\times 512$ pixels. Specifically, 2,212 slices from 18 scans are used for training, and 1,567 slices from 12 scans are used for testing.
@@ 4.1.7 Kvasir-Seg
The Kvasir-Seg dataset contains 1,000 pathological polyp images and their corresponding annotations. We randomly split the dataset into 880 training images and 120 testing images.
@@ 4.1.8 Brain Tumor Detection
This dataset is specifically designed for detecting brain tumors using computer vision techniques. The train set and test set includes 6,930 images and 1,980 images, respectively.
Table 4: Quantitative results on Params, FLOPs, DSC, and HD of our approach against other state-of-the-art methods for medical image segmentation on Synapse multi-organ segmentation dataset. Only DSC is exclusively used for the evaluation of individual organ. The symbol $\uparrow$ indicates the larger the better. The symbol $\downarrow$ indicates the smaller the better. The best result is in Blod, and the second best is underlined.
@@
BRAU-Net++ [11] 62.63 17.66 82.47 19.07 87.95 69.10 87.13 81.53 94.71 65.17 91.89 82.26
@@
Table 5: Qualitative results of different methods for Medical Image Segmentation on ISIC-2018 Segmentation and CVC-ClinicDB datasets.
@@
TransUNet [7] 77.05 84.97 94.56 84.77 89.85 79.95 86.70 98.25 87.63 87.34 105.32 38.54
Swin-Unet [9] 81.87 87.43 95.44 90.97 91.28 84.85 88.21 98.72 90.52 91.13 27.18 8.09
HiFormer [33] 81.54 88.44 94.68 90.23 90.82 85.74 90.48 98.62 91.56 90.37 25.51 8.05
BRAU-Net++ [11] 84.01 90.10 95.61 91.18 92.24 88.17 92.94 98.83 93.84 93.06 62.63 23.03

### v2

# colorlinks=true, linkcolor=blue, anchorcolor=blue, citecolor=blue

@@ Abstract
[…]
@@ 1 Introduction
Deep learning has been widely applied in biomedical research [1], [2], particularly in medical image recognition tasks, such as classification, segmentation, and detection, which are extremely important for computer-aided diagnosis due to their ability to rapidly identify lesion types and accurately localize affected regions [3], [4]. Recently, vision transformer-based approaches demonstrate their effectiveness in addressing these medical recognition tasks. However, due to the diversity and complexity of medical images across different imaging modalities (e.g., MRI, CT, and Optics) and varying image quality (e.g., signal-to-noise ratio and artifacts), most existing vision transformer architectures are typically tailored for task-specific scenarios, such as disease diagnosis [5], [6], [7], [8], anatomical structure and lesion segmentation [9], [10], [11], [12], [13], and abnormality detection [14], [15]. While effective in their respective applications, such task-specific designs may lead to unsatisfactory performance when applied to diverse medical image recognition tasks. This limitation underscores the need for a general-purpose transformer architecture with task-agnostic, modality-invariant, and quality-adaptive characteristics.
@@
The main technical contributions—hierarchical feature representations through a pyramid structure and the DSSA mechanism—endow MedFormer with a strong ability to effectively model long-range dependencies. As a result, the proposed MedFormer architecture serves as a robust general-purpose backbone without any additional bells and whistles. This makes it particularly appropriate for medical recognition tasks such as disease diagnosis, anatomical structure segmentation, and abnormality detection. To the best of our knowledge, this is the first work to attempt developing a general-purpose vision transformer backbone for the medical image recognition community. Overall, our main contributions can be summarized as follows:
@@
[…]
@@ 3.1.3 Matrix Multiplication
To enable matrix multiplication between $\mathbf{A}^{P}$ and $\mathbf{V}^{gg}$, we introduce an additional dimension to the third axis of $\mathbf{A}^{P}$, resulting in a tensor shape of $S^{2}\times\frac{HW}{S^{2}}\times 1\times k_{2}$. Furthermore, to enrich local feature information, we incorporate a local context enhancement module, denoted as $\mathrm{LCE}(\cdot)$, which is constructed using a $\times5$ depth-wise convolution. This module further strengthens feature representation. The final output of the model is expressed as:
@@
Table 1: Variants of MedFormer for medical image classification. $C_{i}$: Channel dimension at stage ii. $S$: the number of regions in feature maps. $k_{1}$: Select the $k_{1}$ most relevant regions. $\lambda$: proportionality coefficient used for the second sparse selection.
@@ 3.2 Complexity Analysis of DSSA
$\mathrm{FLOPs}_{proj}=\underbrace{3HWC^{2}}_{\mathrm{projection}}$
$\mathrm{FLOPs}_{region}=\underbrace{S^{2}(S^{2}C)}_{\mathbf{Q}^{r}\cdot(\mathbf{K}^{r})^{T}}+\underbrace{S^{2}(S^{2}+k_{1})}_{\mathrm{Topk_{1}}}\\$
$\mathrm{FLOPs}_{token}=\underbrace{HW\frac{{k_{1}}HW}{S^{2}}C}_{\mathbf{Q}\cdot(\mathbf{K}^{g})^{T}}+\underbrace{HW(\frac{{k_{1}}HW}{S^{2}}+k_{2})}_{\mathrm{Topk_{2}}}+\underbrace{HW{k_{2}}C}_{\mathbf{A}^{P}\cdot\mathbf{V}^{gg}}$
@@
$=3HWC^{2}+S^{2}(S^{2}C+S^{2}+k_{1})$
$+HW(C+1)(\frac{{k_{1}}HW}{S^{2}}+k_{2}),$
where $S^{2}C>S^{2}+k_{1}$ and $\frac{{k_{1}}HW}{S^{2}}>k_{2}$, we then substitute $S^{2}C$ and $\frac{{k_{1}}HW}{S^{2}}$ into Eq. 11 to obtain the inequality:
$\mathrm{FLOPs}<3HWC^{2}+2(S^{2})^{2}C+\frac{2k_{1}(HW)^{2}}{S^{2}}(C+1).$ (12)
Next, we denote the right hand side of the inequality in Eq. 12 as $\mathrm{InEq}$. Applying the AM-GM inequality $a+b+c\geq 3(abc)^{\frac{1}{3}}$, we proceed with further analysis:
$\mathrm{InEq}$ $=3HWC^{2}+2(S^{2})^{2}C+\frac{2k_{1}(HW)^{2}}{S^{2}}(C+1)$ (13)
$>3HWC^{2}+2(S^{2})^{2}C+\frac{2k_{1}(HW)^{2}}{S^{2}}C$
$=3HWC^{2}+2(S^{2})^{2}C+\frac{k_{1}(HW)^{2}}{S^{2}}C+\frac{k_{1}(HW)^{2}}{S^{2}}C$
$\geq 3HWC^{2}+3\left(2(S^{2})^{2}C\cdot\frac{k_{1}(HW)^{2}}{S^{2}}C\cdot\frac{k_{1}(HW)^{2}}{S^{2}}C\right)^{\frac{1}{3}}$
$=3HWC^{2}+6C(k_{1})^{\frac{2}{3}}(HW)^{\frac{4}{3}}.$
Combining Eq. 12 and Eq. 13, it is evident that the upper bound of FLOPs is strictly less than the lower bound of $\mathrm{InEq}$:
$\mathrm{FLOPs}<3HWC^{2}+6C(k_{1})^{\frac{2}{3}}(HW)^{\frac{4}{3}}.$ (14)
@@
Classification ISIC-2018 Classification[38][39] $\times224$ 7 9376 - 2344
ColonPath [40] $\times224$ 2 8007 - 2002
Brain Tumor [41] $\times224$ 4 5618 - 1405
Segmentation ISIC-2018 Segmentation [38][39] $\times224$ 2 1868 467 259
CVC-ClinicDB [42] $\times256$ 2 441 110 61
Synapse [43] $\times224$ 9 2212 - 1567
Detection Kvasir-Seg [44] $\times256$ 2 800 - 200
Brain Tumor Detection [45] $\times256$ 4 7128 - 1782
Table 3: Comparison of different methods on three medical image classification datasets. * indicates that the model is initialized with weights pretrained on the ImageNet dataset.
Methods Params (M) FLOPs (G) Throughput (img/s) Mem (MB) Top-1 Acc (%)
@@
HiFuse-Tiny[29] 120.00 17.56 415.6 5599 84.61 98.70 95.14
MIL-VT[23] 21.80 4.28 1288.0 1027 84.21 97.55 87.13
KAT [32] 20.08 3.01 879.8 1975 88.54 98.22 96.31
MG-Trans [33] 22.12 3.38 926.7 2033 88.32 98.46 97.16
ViT-S[16] 21.67 4.61 1299.5 1027 87.75 97.30 88.21
PVT-S[21] 23.98 3.83 917.0 1985 88.76 97.30 90.52
Swin-T[17] 27.53 4.51 802.8 2323 87.72 98.10 91.43
BiFormer-S[18] 25.03 4.49 996.6 2091 86.22 98.35 96.33
MedFormer-T 3.20 0.58 1569.3 1715 85.51 97.29 91.57
MedFormer-S 12.63 2.18 941.2 2215 88.24 98.10 95.62
MedFormer-B 28.30 4.79 600.3 2993 89.71 98.72 98.13
MedFormer-T* 3.20 0.58 1569.3 1715 93.61 100.00 99.64
MedFormer-S* 12.63 2.18 941.2 2215 94.44 100.00 99.83
MedFormer-B* 28.30 4.79 600.3 2993 94.57 100.00 99.91
@@ 3.3.1 MedFormer Block
Utilizing DSSA, we construct a basic block called the MedFormer Block. This block is composed of three main components: a $\times3$ depth-wise convolution to implicitly encode relative positional information, a DSSA layer performs sparse selection both at region-level and pixel-level to capture critical features, and a two-layer MLP module with expansion ratio ee=3. Each component incorporates residual connections to maintain the model’s learning capacity, and layer normalization is applied to stabilize the training process and improve the capability of generalization. The detailed design of the block is shown in Fig. 3.
@@ 3.3.2 Architecture
To produce hierarchical feature maps for enhancing performance in classification and various dense prediction tasks, we construct a hierarchical pyramid medical backbone network based on MedFormer Blocks, named MedFormer. The detailed design is illustrated in Fig. 3. Specifically, Stage 1 consists of a patch embedding layer and $N_{1}$ MedFormer Blocks. The patch embedding layer is composed of two $\times3$ convolutions, which transform the input image into a feature map with a resolution of $\frac{H}{4}\times\frac{W}{4}\times C$. To obtain a hierarchical representation, Stage 2 introduces a patch merging layer with a $\times3$ convolution, which reduces the number of tokens and sets the channel dimension to $2C$, resulting in a feature map size of $\frac{H}{8}\times\frac{W}{8}\times 2C$. Then, $N_{2}$ MedFormer Blocks are applied to extract effective features. After repeating the process twice, Stage 3 and Stage 4 generate feature maps with dimensions $\frac{H}{16}\times\frac{W}{16}\times 4C$ and $\frac{H}{32}\times\frac{W}{32}\times 8C$, respectively. Additionally, by adjusting the parameters of MedFormer, we proposed three different variants: Tiny, Small and Base. Table 1 presents the detailed parameter configurations of these MedFormer variants for classification tasks.
@@ 4.1 Datasets
This subsection introduces the datasets used for each task, with detailed information about each dataset provided in Table 2. Notably, for classification and detection datasets, we split them into training and test sets in an 8:2 ratio. As for segmentation datasets, we follow the same data partitioning scheme as previous works, such as Swin-Unet, MISSFormer, and BRAU-Net++.
@@ 4.1.1 ISIC-2018 Classification
[…]
@@ 4.1.2 ColonPath
This is a binary classification dataset from Ruijin Hospital in China, designed to detect malignant cell regions in tissue slides. These images are derived from colonoscopy results of 396 patients. We randomly split the dataset into 8,007 training images and 2,002 test images.
@@ 4.1.3 Brain Tumor
This is an MRI image dataset for four-class brain tumor classification, consisting of a total of 7,023 images, including 5,618 training images and 1,405 test images.
@@ 4.1.4 ISIC-2018 Segmentation
This segmentation dataset is collocated from the ISIC-2018 Challenge Task 1 and contains 2,594 images with varying resolutions. Following [13], we first allocate 10% as the test set, then split the remaining images into training and validation sets with 8:2 ratio.
@@ 4.1.5 CVC-ClinicDB
CVC-ClinicDB is a widely used polyp segmentation dataset consisting of 612 images. Following the same division method as the ISIC-2018 Segmentation dataset, we obtain 441 training images, 110 validation images, and 61 test images.
@@ 4.1.6 Synapse
This is a multi-organ segmentation dataset consists of 30 abdominal CT scans, with each CT volume containing between 85 and 198 slices of $\times512$ pixels. The dataset is partitioned following [11], [37]. Specifically, 2,212 slices from 18 scans are used for training, and 1,567 slices from 12 scans are used for testing.
@@ 4.1.7 Kvasir-Seg
The Kvasir-Seg dataset contains 1,000 pathological polyp images and their corresponding annotations. We randomly split the dataset into 800 training images and 200 testing images.
@@ 4.1.8 Brain Tumor Detection
This dataset is specifically designed for detecting brain tumors using computer vision techniques. The train set and test set includes 7,128 images and 1,782 images, respectively.
Table 4: Quantitative results on Params, FLOPs, DSC, and HD of our approach against other state-of-the-art methods for medical image segmentation on Synapse multi-organ segmentation dataset. Only DSC is exclusively used for the evaluation of individual organ. The symbol $\uparrow$ indicates the larger the better. The symbol $\downarrow$ indicates the smaller the better. The best result is in Blod, and the second best is underlined. * indicates the model is initialized with pre-trained weights.
@@
MedFormer-B 31.28 7.49 81.75 18.72 87.62 71.66 82.25 81.04 95.14 62.12 92.52 81.65
TransUNet* [9] 105.28 29.35 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
Swin-Unet* [11] 27.17 6.16 79.13 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
HiFormer* [35] 25.51 8.05 80.39 14.70 86.21 65.69 85.23 79.77 94.61 59.52 90.99 81.08
BRAU-Net++* [13] 62.63 17.66 82.47 19.07 87.95 69.10 87.13 81.53 94.71 65.17 91.89 82.26
@@
Table 5: Quantitative results of different methods for medical image segmentation on ISIC-2018 Segmentation and CVC-ClinicDB datasets. * indicates the model is initialized with pre-trained weights.
@@
MISSFormer [37] 81.45 88.29 94.81 91.37 89.13 83.68 89.50 98.58 91.65 88.08 42.53 11.92
MedFormer-B 81.67 88.42 94.78 91.15 89.60 84.49 89.99 98.71 91.13 89.55 29.38 7.15
TransUNet* [9] 77.05 84.97 94.56 84.77 89.85 79.95 86.70 98.25 87.63 87.34 105.32 38.54
Swin-Unet* [11] 81.87 87.43 95.44 90.97 91.28 84.85 88.21 98.72 90.52 91.13 27.18 8.09
HiFormer* [35] 81.54 88.44 94.68 90.23 90.82 85.74 90.48 98.62 91.56 90.37 25.51 8.05
BRAU-Net++* [13] 84.01 90.10 95.61 91.18 92.24 88.17 92.94 98.83 93.84 93.06 62.63 23.03